# Principled Foundations for Preference $**{}^{\mbox{\small*}}$

### v1

@@ 2.2 Stochastic choice with abstention
Putting a density on $\alpha$ yields a sampling process totally compatible with the triplet sampling process $(x,y,y^{\prime})$ against which preference optimization is usually done. If deemed preferable or necessary (if e.g. picking $\alpha$ does not follow a stochastic process), one could also consider having the user face lotteries directly in the choices to make. Our framework can still work as is provided one makes the additional "Reduction of Compound Lotteries Axiom" (Machina and Siniscalchi, 2014). Fix $\alpha\in[0,1]$. Let $\mathcal{Y}^{\alpha}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\{(yy^{\prime})_{\alpha}:y,y^{\prime}\in\mathcal{Y}\}$ be the set of available lotteries. An important point to remember is that lottery $(yy)_{\alpha}$ is equivalent to alternative $y$ because any choice probability involving $(yy)_{\alpha}$ breaks down to the one involving only $y$ via (4). Hence $\mathcal{Y}^{\alpha}$contains" $\mathcal{Y}$. We say that choice probabilities are expandable to denote that the support of all actions $\mathcal{Y}$ is extended to cover all lotteries.
@@
The large number of parameters may create an impression of complexity which hides a simple core message: condition (iii) may be read "${\color[rgb]{1,0,0}L_{1}}$ is preferred to ${\color[rgb]{1,.5,0}L_{2}}$ more than ${\color[rgb]{0,0.8046875,0}L_{4}}$ is preferred to ${\color[rgb]{0,1,1}L_{5}}$" and condition (iv) is "${\color[rgb]{1,.5,0}L_{2}}$ is preferred to ${\color[rgb]{1.0,0.88,0.21}L_{3}}$ more than ${\color[rgb]{0,1,1}L_{5}}$ is preferred to ${\color[rgb]{.5,0,.5}L_{6}}$". The right consequence in (7) is then intuitive: "${\color[rgb]{1,0,0}L_{1}}$ is preferred to ${\color[rgb]{1.0,0.88,0.21}L_{3}}$ more than ${\color[rgb]{0,0.8046875,0}L_{4}}$ is preferred to ${\color[rgb]{.5,0,.5}L_{6}}$". The left consequence in (7) imposes that the wedge on condition (ii) be in fact a triangle in $\mathcal{E}_{x}^{\alpha}$, imposing no abstention when choosing ${\color[rgb]{0,0.8046875,0}L_{4}}$ vs ${\color[rgb]{.5,0,.5}L_{6}}$ (which is natural considering (i) and (ii)). Importantly, monotonicity needs to be defined only for one non-trivial $\alpha$ ($\in(0,1)$), which is different from the requirements of a LCS.
@@ 3.2 Generalization of step (1)
$J_{Q,r}(\pi(.|x))\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\mathbb{E}_{y\sim\pi(.|x)}[r(x,y)]-Q(\pi(.|x)\|\pi_{\mbox{\tiny{ref}}}(.|x))$ (11)
Such a general formulation is ubiquitous in ML. The question we address is: what form should have $Q$ There is a simple condition that imposes a simple analytical structure generalizing the KL divergence. Suppose we have learned
$\pi_{*}(.|x)$ $\stackrel{{\scriptstyle\mathrm{.}}}{{=}}$ $\arg\max J_{Q,r}(\pi(.|x)),$
@@
$Q(\pi(.|x)\|\pi^{\prime}(.|x))$ $=$ $K+D_{\phi,G}(\pi(.|x)\|\pi^{\prime}(.|x)),$ (13)
where $K$ does not depend on $\pi(.|x)$ and $D_{\phi,G}$ is a Bregman divergence. More specifically, we can show $\phi(\bm{p})=-L(\bm{p},\bm{p})$ for some proper loss $\ell$ (1) and can pick $G_{i}=-\ell_{i}$.
Proof in Appendix, Section VII.1. Note that $G_{i}$ is allowed to depend not just on $y_{i}$.
@@ 3.3 Generalization of step (2)
It is well known since Amari and Nagaoka (2000) that the solution to (11) with $Q$ a Bregman divergence introduces a canonical mapping between policy and rewards via the selection $\bm{G}$, called a dually flat coordinate system. In ML, it has such a broad utility that $\bm{G}$ has a name: it is a [^2] If $\phi$ is differentiable, it is ”the” canonical link, unique. canonical link of the loss $\bm{\ell}$ (Reid and Williamson, 2010). Step (2) in DPO relies on a similar mapping but the policy is replaced with the choice probabilities. Lacking an obvious function to make the link as would be in Step (1), we have to resort to the axiomatization of $p(y\succ y^{\prime}|x)$ to create this mapping. It is done by the following Theorem.
@@
such that $p(y\succ y^{\prime}|x)=F(u(x,y)-u(x,y^{\prime}))$.
@@
Here, $\lambda$ is the Lagrange multiplier enforcing normalization of $\pi(.|x)$, $\mu_{y_{i}}\geq 0$ is the one for the constraint of non-negativity for $\pi(y_{i}|x)$; finally, for a predicate $\Pi$, $0_{\Pi}$ takes value 0 if $\Pi$ is true and 1 otherwise. If we plug $u(x,y_{i})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}r(x,y_{i})+0_{\pi(y_{i}|x)>0}\cdot\mu_{y_{i}}$ (if all policy values are $>0$ such as in DPO, this is just the reward function) in (14), then we remark for any $x,y,y^{\prime}$,
@@ 3.4 Generalization of step (3) with a canonical connection
  • this problem is equivalent to a binary classification problem where $(x,y_{i},y_{j})$ is positive if say "$y_{i}\succ y_{j}$ given $x$" and negative if "$y_{j}\succ y_{i}$ given $x$" (Sun et al., 2025) so the proper loss is scalar valued, which we note $L(p,p)=p\ell_{1}(p)+(1-p)\ell_{0}(p)$ after Reid and Williamson (2010). The indexing in $\{0,1\}$ is a convention departing from our indexing in $1,2,...,n$ (Definition 1): "0" usually refers to the so-called "negative class" and "1" to the "positive class". Such a loss is noted $(\ell_{0},\ell_{1})$;
  • most DPO generalizations have so far considered a restriction to symmetric losses (Nock and Nielsen, 2008), for which we have $\ell_{0}(p)=\ell_{1}(1-p)$ and (16) implicitly assumes the usual convention =.$\ell\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\ell_{1}$.
We now tackle the problem from the most general standpoint and drop the symmetry assumption. What are conditions for which (16) bears relevant properties in the context of Savage’s theory, given that the arguments of "$\ell\circ F$" in (16) are reals and not probabilities ?
@@
For any strictly proper loss $(\ell_{0},\ell_{1})$, suppose we pick ${\color[rgb]{1,0,0}F=H^{-1}}$ in Theorem 6 where $H$ is as in Theorem 7 and let =.$\ell\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\ell_{0}$ in (16). Then (16) simplifies to:
@@
(the modelling "price" for never abstaining) Forbidding abstention yields to stochastic choice models pioneered by Debreu, enforcing $F(z)+F(-z)=1$. Here comes an interesting fact: the simplification using the convex conjugate in Corollary 8 and commonly used in the state of the art DPO-style analysis happens only if the strictly proper loss chosen in Theorem 7 is symmetric.
@@ 3.5 Further generalization of step (3) with a composite connection
In the canonical connection used in Corollary 8, a strictly proper loss $(\ell_{0},\ell_{1})$ defines both $\ell$ and $F$ in the objective $I(\pi_{\bm{\theta}})$ (16), $\ell$ being $\ell_{0}$ and $F$ the inverse of a canonical link of the loss. This yields a broad family of losses that rely on the convex surrogate of the loss, $\phi^{\star}(-z)$, but as we said, they all share the same blueprint (1-Lipschitz, etc.). Let us strive for the maximum generality in covering "acceptable" losses for the objective $I(\pi_{\bm{\theta}})$ (16): what is the set of functions $\psi:\mathbb{R}\rightarrow\mathbb{R}$ that could be used in lieu of $\phi^{\star}$ while still guaranteeing the existence of (i) a choice function $F$ meeting the conditions of Theorem 6, (ii) a strictly proper loss $(\ell_{0},\ell_{1})$, and such that $\psi=\ell_{0}\circ F$ (we refer to such a connection as being "composite"). The result, whose proof is constructive, shows a very broad result. First, Lipschitzness is (predictably) optional, but convexity is also optional. Second, for any applicable $\psi$, any $F$ meeting the conditions of Theorem 6 can work. So any applicable $\psi$ can be operated with any KLST model.
@@
For any strictly increasing function $\psi:\mathbb{R}\rightarrow\mathbb{R}$, any strictly increasing function $F:\mathbb{R}\rightarrow[0,1]$ satisfying $\mathrm{dom}F\subseteq\mathrm{dom}\psi$, there exists a strictly proper loss $(\ell_{0},\ell_{1})$ such that
@@
Note that we cannot have a result in which we would instead pick any strictly proper loss and then get $F$ meeting the conditions of Theorem 6. Here is a simple counterexample: pick $\psi(z)\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\exp(z)$ (gives the exponential loss), and $\ell_{0}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}-\log(1-p)$ (the log-loss). We get $F=\ell_{0}^{-1}\circ\psi=1-\exp(-\exp(z))$, which violates the condition $F(-z)+F(z)\leq 1,\forall z\in\mathbb{R}$ ($F(0)=1-(1/e)>1/2$).
@@ 4 Discussion
{mdframed}
[style=MyFrame]
          pmpo     pppo     $\phi$-po
      Ingredients     $\psi\nearrow\nearrow$     $(\ell^{a}_{0},\ell^{a}_{1})$ strictly proper     $\phi$ eligible
        $\bm{\ell}^{b}$ proper     Def. 14
      $\psi$ in (17)     $\psi$     $\phi^{\star}$ (Thm. 7)     Thm. 15
        ($\ell^{a}_{0}$, $F$: Thm. 6)     ($F=(\ell^{a}_{0}(p)-\ell^{a}_{1}(p))^{-1}$)     & Def. 17
      $G_{i}$ in (17)     $\ell^{b}_{i}$     Def. 17
@@ 4.1 Proper-Monotone Preference Optimization (pmpo)
For any strictly increasing $\psi:\mathbb{R}\rightarrow\mathbb{R}$ and $\bm{\ell}^{b}$ (proper), the pmpo Preference Optimization framework is the setting of the following parameters: =.bsuperscript.$\bm{\ell}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\bm{\ell}^{b}$ in Theorem 5 and $\psi$ used as is in (17).
@@ 4.2 Proper-Proper Preference Optimization (pppo)
For any losses $(\ell^{a}_{0},\ell^{a}_{1})$ (strictly proper) and $\bm{\ell}^{b}$ (proper), the pppo Preference Optimization framework is the setting of the following parameters in Theorems 5, 6 and 7: (i) =.bsuperscript.$\bm{\ell}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\bm{\ell}^{b}$ in Theorem 5, (ii) $F\stackrel{{\scriptstyle\mathrm{.}}}{{=}}(\ell^{a}_{0}(p)-\ell^{a}_{1}(p))^{-1}$ in Theorem 6, (iii) $(\ell_{0},\ell_{1})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}(\ell^{a}_{0},\ell^{a}_{1})$ in Theorem 7.
@@ 4.3 $\phi$-Preference Optimization ($\phi$-po)
The $\phi$-po Preference Optimization framework is pppo with $(\ell^{a}_{0},\ell^{a}_{1})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}(\ell_{0},\ell_{1})$ as per (19) and $\bm{\ell}^{b}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}[\ell^{b}_{i}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\tilde{\ell}_{1}]_{i=1}^{n}$, where ~$\tilde{\ell}_{1}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}(-\phi(p)-\phi(1-p)-(1-p)(H(p)-H(1-p)))/2$.
Note that ~$\tilde{\ell}_{1}$ symmetrizes the loss as remarked in Theorem 15. When $\phi$ is differentiable, we immediately get $\phi^{\star}(z)=z\phi^{\prime-1}(z)-\phi(\phi^{\prime-1}(z))$ for the convex conjugate. Under additional assumptions on $\phi$, functions $G_{i}$ in Theorem 5 adopt a remarkable analytical form.

### v2

# Wenxuan Zhou Google DeepMind Shujian Zhang Google DeepMind Brice Magdalou CEE-M, Montpellier U. John Lambert Google DeepMind Ehsan Amid Google DeepMind Richard Nock Google Research Andrew Hard Google DeepMind

@@ 2.2 Stochastic choice with abstention
Putting a density on $\alpha$ yields a sampling process totally compatible with the triplet sampling process $(x,y,y^{\prime})$ against which preference optimization is usually done. If deemed preferable or necessary (if e.g. picking $\alpha$ does not follow a stochastic process), one could also consider having the user face lotteries directly in the choices to make. Our framework can still work as is provided one makes the additional "Reduction of Compound Lotteries Axiom" (Machina and Siniscalchi, 2014). Fix $\alpha\in[0,1]$. Let $\mathcal{Y}^{\alpha}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\{(yy^{\prime})_{\alpha}:y,y^{\prime}\in\mathcal{Y}\}$ be the set of available lotteries. An important point to remember is that lottery $(yy)_{\alpha}$ is equivalent to alternative $y$ because any choice probability involving $(yy)_{\alpha}$ breaks down to the one involving only $y$ via (4). Hence $\mathcal{Y}^{\alpha}$ "contains" $\mathcal{Y}$. We say that choice probabilities are expandable to denote that the support of all actions $\mathcal{Y}$ is extended to cover all lotteries.
@@
The large number of parameters may create an impression of complexity which hides a simple core message: condition (iii) may be read $\color[rgb]{1,0,0}L_{1}$} is preferred to $\color[rgb]{1,.5,0}L_{2}$} more than $\color[rgb]{0,0.8046875,0}L_{4}$} is preferred to $\color[rgb]{0,1,1}L_{5}$}" and condition (iv) is $\color[rgb]{1,.5,0}L_{2}$} is preferred to $\color[rgb]{1.0,0.88,0.21}L_{3}$} more than $\color[rgb]{0,1,1}L_{5}$} is preferred to $\color[rgb]{.5,0,.5}L_{6}$}". The right consequence in (7) is then intuitive: $\color[rgb]{1,0,0}L_{1}$} is preferred to $\color[rgb]{1.0,0.88,0.21}L_{3}$} more than $\color[rgb]{0,0.8046875,0}L_{4}$} is preferred to $\color[rgb]{.5,0,.5}L_{6}$}". The left consequence in (7) imposes that the wedge on condition (ii) be in fact a triangle in $\mathcal{E}_{x}^{\alpha}$, imposing no abstention when choosing $\color[rgb]{0,0.8046875,0}L_{4}$} vs $\color[rgb]{.5,0,.5}L_{6}$} (which is natural considering (i) and (ii)). Importantly, monotonicity needs to be defined only for one non-trivial $\alpha$ ($\in(0,1)$), which is different from the requirements of a LCS.
@@ 3.2 Generalization of step (1)
$J_{R,r}(\pi(.|x))\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\mathbb{E}_{y\sim\pi(.|x)}[r(x,y)]-R(\pi(.|x)\|\pi_{\mbox{\tiny{ref}}}(.|x)),$ (11)
where $R$ is the regret of a function $Q(\pi(.|x)\|\pi_{\mbox{\tiny{ref}}}(.|x))$ that we aim to elicit. The regret being the difference between $Q$ and its infimal value (assumed finite wlog), it is non-negative. Such a general formulation is ubiquitous in ML. The question we address is: what form should have $Q$ ? There is a simple condition that imposes a simple analytical structure generalizing the KL divergence. Suppose we have learned
$\pi_{*}(.|x)$ $\stackrel{{\scriptstyle\mathrm{.}}}{{=}}$ $\arg\max J_{R,r}(\pi(.|x)),$
@@
$Q(\pi(.|x)\|\pi^{\prime}(.|x))$ $=$ $K(\pi(.|x))+D_{\phi,G}(\pi(.|x)\|\pi^{\prime}(.|x)),$ (13)
where $D_{\phi,G}$ is a Bregman divergence. More specifically, we can show $\phi(\bm{p})=-L(\bm{p},\bm{p})$ for some proper loss $\ell$ (1) and can pick $G_{i}=-\ell_{i}$.
Proof in Appendix, Section VII.1. Note that $G_{i}$ is allowed to depend not just on $y_{i}$. It comes from Williamson et al. (2016, Proposition 7) that $Q(\pi(.|x)\|\pi(.|x))=K(\pi(.|x))$, so we get the expression of the regret in (11),
$R(\pi(.|x)\|\pi^{\prime}(.|x))$ $\stackrel{{\scriptstyle\mathrm{.}}}{{=}}$ $Q(\pi(.|x)\|\pi^{\prime}(.|x))-Q(\pi(.|x)\|\pi(.|x))$
$=$ $D_{\phi,G}(\pi(.|x)\|\pi^{\prime}(.|x)).$
@@ 3.3 Generalization of step (2)
[…]
@@
such that
$p(y\succ y^{\prime}|x)$ $=$ $F(u(x,y)-u(x,y^{\prime})).$ (14)
@@
Here, $\lambda$ is the Lagrange multiplier enforcing normalization of $\pi(.|x)$, $\mu_{y_{i}}\geq 0$ is the one for the constraint of non-negativity for $\pi(y_{i}|x)$; finally, for a predicate $\Pi$, $0_{\Pi}$ takes value 0 if $\Pi$ is true and 1 otherwise. If we plug $u(x,y_{i})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}r(x,y_{i})+0_{\pi(y_{i}|x)>0}\cdot\mu_{y_{i}}$ (if all policy values are >0>0 such as in DPO, this is just the reward function) in (15), then we remark for any $x,y,y^{\prime}$,
@@ 3.4 Generalization of step (3) with a canonical connection
  • this problem is equivalent to a binary classification problem where $(x,y_{i},y_{j})$ is positive if say $y_{i}\succ y_{j}$ given xx" and negative if $y_{j}\succ y_{i}$ given xx" (Sun et al., 2025) so the proper loss is scalar valued, which we note $L(p,p)=p\ell_{1}(p)+(1-p)\ell_{0}(p)$ after Reid and Williamson (2010). The indexing in $\{0,1\}$ is a convention departing from our indexing in 1,2,,n1,2,...,n (Definition 1): "0" usually refers to the so-called "negative class" and "1" to the "positive class". Such a loss is noted $(\ell_{0},\ell_{1})$;
  • most DPO generalizations have so far considered a restriction to symmetric losses (Nock and Nielsen, 2008), for which we have $\ell_{0}(p)=\ell_{1}(1-p)$ and (17) implicitly assumes the usual convention $\ell\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\ell_{1}$.
We now tackle the problem from the most general standpoint and drop the symmetry assumption. What are conditions for which (17) bears relevant properties in the context of Savage’s theory, given that the arguments of $\ell\circ F$" in (17) are reals and not probabilities ?
@@
For any strictly proper loss $(\ell_{0},\ell_{1})$, suppose we pick $\color[rgb]{1,0,0}F=H^{-1}$} in Theorem 6 where $H$ is as in Theorem 7 and let $\ell\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\ell_{0}$ in (17). Then (17) simplifies to:
@@
(the modelling "price" for never abstaining) Forbidding abstention yields to stochastic choice models pioneered by Debreu, enforcing F(z)+F(z)=1F(z)+F(-z)=1. Here comes an interesting fact: the simplification using the convex conjugate in Corollary 8 and commonly used in the state of the art DPO-style analysis happens only if the strictly proper loss chosen in Theorem 7 is symmetric.
@@ 3.5 Further generalization of step (3) with a composite connection
In the canonical connection used in Corollary 8, a strictly proper loss $(\ell_{0},\ell_{1})$ defines both $\ell$ and $F$ in the objective $I(\pi_{\bm{\theta}})$ (17), $\ell$ being $\ell_{0}$ and $F$ the inverse of a canonical link of the loss. This yields a broad family of losses that rely on the convex surrogate of the loss, $\phi^{\star}(-z)$, but as we said, they all share the same blueprint (1-Lipschitz, etc.). Let us strive for the maximum generality in covering "acceptable" losses for the objective $I(\pi_{\bm{\theta}})$ (17): what is the set of functions $\psi:\mathbb{R}\rightarrow\mathbb{R}$ that could be used in lieu of $\phi^{\star}$ while still guaranteeing the existence of (i) a choice function $F$ meeting the conditions of Theorem 6, (ii) a strictly proper loss $(\ell_{0},\ell_{1})$, and such that $\psi=\ell_{0}\circ F$ ? (we refer to such a connection as being "composite"). The result, whose proof is constructive, shows a very broad result. First, Lipschitzness is (predictably) optional, but convexity is also optional. Second, for any applicable $\psi$, any $F$ meeting the conditions of Theorem 6 can work.
@@
For any strictly increasing function $\psi:\mathbb{R}\rightarrow\mathbb{R}$, any strictly increasing function $F:\mathbb{R}\rightarrow[0,1]$, there exists a strictly proper loss $(\ell_{0},\ell_{1})$ such that
@@
Note that we cannot have a result in which we would instead pick any strictly proper loss and then get $F$ meeting the conditions of Theorem 6. Here is a simple counterexample: pick $\psi(z)\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\exp(z)$ (gives the exponential loss), and $\ell_{0}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}-\log(1-p)$ (the log-loss). We get $F=\ell_{0}^{-1}\circ\psi=1-\exp(-\exp(z))$, which violates the condition $F(-z)+F(z)\leq 1,\forall z\in\mathbb{R}$ (F(0)=1(1/e)>1/2F(0)=1-(1/e)>1/2).
Consequences of Theorem 11
Theorem 11 does not just widen the scope of our framework: it hides a crucial message for the experimenter and experimental design in a field where variations to DPO are numerous. Experiments would start by the choice that defines the loss, $\psi$ (18), from which embedding it in our framework would necessitate to find a working model of social choice whose choice probabilities have a KLST structure. Given a reward / utility function $u$ (implicit), Theorem 11 says that this second step is not necessary: any $\psi$ applicable according to Theorem 11 can operate with any model whose choice probabilities have a KLST structure, the hidden "link" operating the canonical or composite connection being a Savage loss.
@@ 4 Discussion
          pmpo     pppo     $\phi$-po           Ingredients     $\psi\nearrow\nearrow$     $(\ell^{a}_{0},\ell^{a}_{1})$ strictly proper     $\phi$ eligible             $\bm{\ell}^{b}$ proper     Def. 14           $\psi$ in (18)     $\psi$     $\phi^{\star}$ (Thm. 7)     Thm. 15             ($\ell^{a}_{0}$, $F$: Thm. 6)     ($F=(\ell^{a}_{0}(p)-\ell^{a}_{1}(p))^{-1}$)     & Def. 17           $G_{i}$ in (18)     $\ell^{b}_{i}$     Def. 17
@@ 4.1 Proper-Monotone Preference Optimization (pmpo)
For any strictly increasing $\psi:\mathbb{R}\rightarrow\mathbb{R}$ and $\bm{\ell}^{b}$ (proper), the pmpo Preference Optimization framework is the setting of the following parameters: $\bm{\ell}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\bm{\ell}^{b}$ in Theorem 5 and $\psi$ used as is in (18).
@@ 4.2 Proper-Proper Preference Optimization (pppo)
For any losses $(\ell^{a}_{0},\ell^{a}_{1})$ (strictly proper) and $\bm{\ell}^{b}$ (proper), the pppo Preference Optimization framework is the setting of the following parameters in Theorems 5, 6 and 7: (i) $\bm{\ell}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\bm{\ell}^{b}$ in Theorem 5, (ii) $F\stackrel{{\scriptstyle\mathrm{.}}}{{=}}(\ell^{a}_{0}(p)-\ell^{a}_{1}(p))^{-1}$ in Theorem 6, (iii) $(\ell_{0},\ell_{1})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}(\ell^{a}_{0},\ell^{a}_{1})$ in Theorem 7.
@@ 4.3 $\phi$-Preference Optimization ($\phi$-po)
The $\phi$-po Preference Optimization framework is pppo with $(\ell^{a}_{0},\ell^{a}_{1})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}(\ell_{0},\ell_{1})$ as per (20) and $\bm{\ell}^{b}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}[\ell^{b}_{i}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\tilde{\ell}_{1}]_{i=1}^{n}$, where $\tilde{\ell}_{1}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}(-\phi(p)-\phi(1-p)-(1-p)(H(p)-H(1-p)))/2$.
Note that $\tilde{\ell}_{1}$ symmetrizes the loss as remarked in Theorem 15. When $\phi$ is differentiable, we immediately get $\phi^{\star}(z)=z\phi^{\prime-1}(z)-\phi(\phi^{\prime-1}(z))$ for the convex conjugate. Under additional assumptions on $\phi$, functions $G_{i}$ in Theorem 5 adopt a remarkable analytical form.